OpenStudy (anonymous):

how do I find the perimeter of a rhombus EHFG if I only know the lengths of EG and FH??? PLEASE HELP!

OpenStudy (anonymous):

Given that EFGH is a rhombus and EG and FH are the two diagonals. We know that diagonals bisect each other at 90 degrees. Let EG and FH meet at O. Consider triangle EFO. OE=EG/2 and OF=FH/2. Now, EF^2=OE^2+OF^2 Or, EF^2=(EG^2/4)+(FH^2/4) (By substitution) Or, EF=[(EG^2/4)+(FH^2/4)]^1/2